OpenStudy (anonymous):

Two lines, C and D, are represented by the equations given below: Line C: y = x + 10 Line D: y = 3x + 2 Which of the following shows the solution to the system of equations and explains why? A. (5,17), because one of the lines passes through this point. B. (5,17), because the point lies between the two axes. C.(4,14), because the point does not lie on any axes . D.(4,14), because both lines pass through the point.

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

Use substitution x+10=3x+2 Solve for x

OpenStudy (christina166):

Hey is this geometry?

pooja195 (pooja195):

It's algebra...

OpenStudy (christina166):

Oh

OpenStudy (anonymous):

3x and x and is the answer 8?

pooja195 (pooja195):

Good now divide by 2 8/2=?

OpenStudy (anonymous):

4

pooja195 (pooja195):

Good now plug that back into equation 2 y=3(4)+2

OpenStudy (anonymous):

14?

OpenStudy (anonymous):

c or d right?

pooja195 (pooja195):

Our x value is is 4 y=14

pooja195 (pooja195):

Which matches the coordinate?

OpenStudy (anonymous):

d?

OpenStudy (anonymous):

thank you!!!!!!!!